OpenStudy (anonymous):

Use the slope formula to find the missing coordinate. Formula: y2-y1 over x2-x1= the slope C (a,b) D(c, ____) m=j I tried using the slope formula but i'm doing something wrong. Please help...

OpenStudy (anonymous):

Does the question tell you anything else? If not: Let the missing coordinate be x. \[x - b \div a - c = j\]x - b = j(a - c) x = j(a - c) + b So D(c, j(a - c)+b)

OpenStudy (anonymous):

Thank u very much for ur help...